What is the concept of 2 to 5 odds and the probability
example
there is a 2 to 5 odds that we will win the lottery, what is the probability that we wont? and that we will?

Answer by edjones(8007)

there is a 2 to 5 odds that we will win the lottery.
That means that the odds in favor of winning are 2 to 5 and the odds against are 5 to 2.
The probability that we will win is 2/(5+2) = 2/7; the probability that we wont is 1 - 2/7 = 5/7.
.
Ed